OpenStudy (anonymous):

A current of 2 A is passed through a coil of resistance 75 ohm for 2 minutes,(a)how much heat energy is produced/ (b)how much charge is passed through the resistance?

OpenStudy (anonymous):

@dan815

OpenStudy (irishboy123):

hint: use \(P (= \frac{dE}{dt}) = I^2R \) and \(I = \frac{dQ}{dt}\)